Bibliography

[A-L-R] M. Aizenman, J.L. Lebowitz, D. Ruelle, Some rigorous results on the Sherrington-Kirkpatrick model, Commun. Math. Phys., 112 (1987) 3–20.

[B-P] A. Bovier, P. Picco (editors), Mathematical Aspects of Spin Glasses and Neural Networks, Progress in Probability, Vol. 41, Birkhäuser, Boston, 1997.

[C] F. Comets, A spherical bound for the Sherrington-Kirkpatrick model, Hommage à P.-A. Meyer et J. Neveu, Astérisque, 236, (1996) 103–108.

[C-N] F. Comets, J. Neveu, The Sherrington-Kirkpatrick model of spin glasses and stochastic calculus: the high temperature case, Comm. Math. Phys., 166, 3 (1995) 549–564.

[D] B. Derrida, Random energy model: An exactly solvable model of disordered systems, Phys. Rev. B, 24, #5 (1981) 2613–2626.

[F-Z] J. Fröhlich, B. Zegarlinski, Some comments on the Sherrington-Kirkpatrick model of spin glasses, Commun. Math. Phys., 112 (1987) 553–566.

[G] E. Gardner, Spin glasses with p-spin interactions, Nuclear Phys. B, 257, #6 (1985) 747–765.

[G-G] S. Ghirlanda, F. Guerra, General properties of overlap probability distributions in disordered spin systems. Towards Parisi ultrametricity, J. Phys. A, 31, #46 (1998) 9149–9155.

[I-S-T] I.A. Ibragimov, V. Sudakov, B.S. Tsirelson, Norms of Gaussian sample functions, “Proceedings of the Third Japan-USSR Symposium on Probability”, Tashkent, 1975, Lecture Notes in Math., 550, Springer Verlag, Berlin, 1976, 20–41.

[K] J.-P. Kahane, Une inegalité du type de Slepian et Gordon sur les processus gaussiens, Israel J. Math., 55, 1 (1986) 109–110.

[K-T-J] J. Kosterlitz, D. Thouless, R. Jones, Spherical model of spin glass, Phys. Rev. Lett., 36 (1976) 1217–1220.

[M-P-V] M. Mézard, G. Parisi, M. Virasoro, Spin glass Theory and beyond, World Scientific, Singapore, 1987.

[P] G. Parisi, Field Theory, Disorder, Simulation, World Scientific Lecture Notes in Physics 45, World Scientific, Singapore, 1992.

106 Bibliography

[Pi] G. Pisier, Probabilistic methods in the geometry of Banach Spaces, Probability and analysis, Varenna 1985, Springer Verlag Lecture Notes in Math. n^{∘} 1206 (1996) 167–241.

[P-Y] J. Pitman, M. Yor, The two-parameter Poisson-Dirichlet distribution derived

from a stable subordinator, Ann. Probab., 25 (1997) 855–900.

[Sh] M. Shcherbina, On the replica-symmetric solution of the Sherrington-Kirkpatrick model, Helv. Phys. Acta, 70 (1997) 838–853.

[S-K] D. Sherrington, S. Kirkpatrick, Solvable model of spin glass, Phys. Rev. Lett.,

35 (1972) 1792–1796.

[T1] M. Talagrand, Concentration of measure and isoperimetric inequalities in product spaces, Publ. Math. I.H.E.S., 81 (1995) 73–205.

[T2] , The Sherrington-Kirkpatrick model: a challenge to mathematicians, Probab. Theory Related Fields, 110 (1998) 109–176.

[T3] , Rigorous low temperature results for the p-spin interaction model, Probab. Theory Related Fields, 117 (2000) 303–360.

[T4] , Exponential inequalities and replica-symmetry breaking for the Sherrington-Kirkpatrick model, Ann. Probab., 28 (2000) 1018–1062.

[T5] , Huge random structures and mean field models for spin glasses, in “Proceedings of the International Congress of Mathematicians, Vol. I (Berlin 1998)”, Documenta Math., Extra Vol. I (1998) 507–536.

[T6] , Verres de spin et optimisation combinatoire, Séminaire Bourbaki, March 1999, Astérisque, 266 (2000) Exp. n^{∘} 859, 287–317.

[T7] , Self organization in a spin glass model, in preparation.

[T8] , On the high temperature region of the Sherrington-Kirkpatrick model, Ann. Probab., to appear.